OpenStudy (anonymous):

Find the mean and the standard deviation of each data set . Round to the nearest hundredth. 1.) 4, 8, 5, 12, 3, 8, 5 and 2 2.) 102, 98, 103, 86, 101, 110 3.) 8.2, 11.6, 8.7, 10.6, 9.4, 10.1, 9.3

OpenStudy (anonymous):

The means are 1.) 41 2.) 600 3.) 67.9

OpenStudy (perl):

are you allowed to use a calculator

OpenStudy (anonymous):

Whats standard deviation?

OpenStudy (anonymous):

yes

OpenStudy (perl):

standard deviation is a measure of variability or 'spread'

OpenStudy (perl):

how spread out the data is relative to the mean

OpenStudy (perl):

$$ \Large st. dev = \sqrt\frac{\sum (x_i - \bar x )^2} {n} $$

OpenStudy (anonymous):

Not understanding..

OpenStudy (perl):

here is a standard deviation calculator https://www.easycalculation.com/statistics/standard-deviation.php